The shortest leg of a right triangle is 2 less than the hypotenuse and the hypotenuse is one longer than the longest leg right in quadratic equation form

Answer by mananth(16949)

longest leg be x
hypotenuse= x+1
shortest leg = x-2
..
x^2+(x-2)^2= (x+1)^2